OpenStudy (anonymous):

Solve the inequality: 4x(x + 1) ≥ 3

OpenStudy (anonymous):

4x^2 +4x-3>=0 (2x-1)(2x+3)>=0 solve it

OpenStudy (anonymous):

will give you: 2x-1>=0 2x+3>=0 first option 2x-1<=0 2x+3<=0 second option

OpenStudy (anonymous):

can you finish it?

OpenStudy (anonymous):

x>=1/2 x<=-3/2

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

i got >1/2 and < -1 1/2

OpenStudy (anonymous):

you are right. only less or equal in both cases - see Safiah answer

OpenStudy (anonymous):

I mean less or equal and more or equal...

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome!